OpenStudy (vera_ewing):

In the galvanic cell in question 1, nickel has a reduction potential of –0.25 V. The total cell potential is +0.59 V. What is the reduction potential of copper? Explain.

OpenStudy (vera_ewing):

@ganeshie8

OpenStudy (vera_ewing):

@Hoslos

OpenStudy (anonymous):

Basically, when you add 2 potentials, you get the resultant one. So make an equation: -0.25+x= +0.59, then x=0.59+0.25=+0.84.

OpenStudy (vera_ewing):

So what is the reduction and explain?

OpenStudy (anonymous):

My mistake. When two metals react like these one, which is Ni2+ .... - Cu, iyou need to consider which one is the more replacing active. In this case is Ni2+. Since Reduction means gain in electrons, then that substance will be negatively charged. Next, knowing that a reaction only takes place when two oppositevily charged particles are together, it means that Ni is negative in potential and Cu is positive. Let us say that the potential for Ni is x and for Cu is y. In an equation it is: -Ni + +Cu = +0.59 Which simplifies to -(-0.25) + Cu=0.59, then the potential for copper is +0.34.

OpenStudy (vera_ewing):

so is that it? that's the answer?

OpenStudy (anonymous):

Yes.

OpenStudy (vera_ewing):

thank you so much! :)))

OpenStudy (anonymous):

My pleasure!